OpenStudy (anonymous):

Surface area and volume of a cubic section using spherical coordinates:

OpenStudy (anonymous):

A section of a sphere is defined with: \[0 \le r \le 2\] \[0 \le \theta \le 90^{o}\] \[30^{o} \le \phi \le 90^{o}\] I need the area and volume of the area described.

OpenStudy (anonymous):

The SA should be around 4.188 and the Volume is 2.79 from using just the formulas.

OpenStudy (anonymous):

I have set up the integrals a dozen times and cant get those numbers. Volume should be something like this, correct? \[\int\limits_{30}^{90}\int\limits_{0}^{90}\int\limits_{0}^{2} R ^{2}\sin \theta d \theta d \phi\]

OpenStudy (anonymous):

And surface area this: \[\int\limits_{0}^{90}\int\limits_{30}^{90} R ^{2} \sin \theta d \theta \]

OpenStudy (anonymous):

@phi - I already owe you a lunch for the last time you helped me out, any chance for an opinion on this one?

OpenStudy (anonymous):

nm I got it :)